A firm must plan production for the next six months. Each unit costs $480 to produce and it has an inventory holding cost of $20 per unit per month based on ending inventory levels. The cost to hire a worker is $330, and the cost to fire a worker is $660 per worker. Each worker produces 10 units per month. There are 25 persons on the payroll at the beginning of the first month. The company currently has 200 units of inventory in stock, and it wants to keep at least this many units as safety stock at the end of every month. Assume hiring and layoff/firing, if necessary, occur at the beginning of the month.

Month 1 2 3 4 5 6
Demand 420 420 420 420 470 550

a. From the information given above, what level production rate will meet demand for the next six months? Using that rate, complete the level plan below. (Leave no cells blank - be certain to enter "0" wherever required.)

Level Plan
Month Demand

Regular

Production

Ending

Inventory

Number of

Workers

Hire Fire
1 420
2 420
3 420
4 420
5 420
6 420
Total 2700

b-1. At that production rate, what is the maximum end-of-period inventory experienced at some time during the six months?

Maximum end-of-period inventory    Units

b-2. What would be the cost of a level production plan?

Total Cost

c-1. From the information given above, Develop a chase (Hire and Fire only) Production Plan? (Leave no cells blank - be certain to enter "0" wherever required.)

Level Plan
Month Demand

Regular

Production

Ending

Inventory

Number of

Workers

Hire Fire
1 420
2 420
3 420
4 420
5 420
6 420
Total 2700

c-2. What is the total cost of a chase (hire and fire only) production plan?

Total Cost
Answer #1

a.

Level Plan
Month Demand Regular Ending Number of Hire Fire
Production Inventory Workers
1 420 450 230 45 20 0
2 420 450 260 45 0 0
3 420 450 290 45 0 0
4 420 450 320 45 0 0
5 470 450 300 45 0 0
6 550 450 200 45 0 0
Total 2700 2700 1600 20 0

b-1.

Maximum end-of-period inventory = 320 units

b-2.

Total cost = regular production cost + inventory holding cost + hire cost + fire cost

= 2700*480+1600*20+20*330+0*660 = 1334600

c-1.

Hire and Fire
Month Demand Regular Ending Number of Hire Fire
Production Inventory Workers
1 420 420 200 42 17 0
2 420 420 200 42 0 0
3 420 420 200 42 0 0
4 420 420 200 42 0 0
5 470 470 200 47 5 0
6 550 550 200 55 8 0
Total 2700 2700 1200 30 0


C-2
Total cost = regular production cost + inventory holding cost + hire cost + fire cost

= 2700*480+1200*20+30*330+0*660 = 1329900